# Multi-type branching processes with immigration generated by point processes

### v1

@@ 5 Proofs
Proof of Theorem 4.3.
We start with the generating function of the process, similarly to (4.3), the generating function of the multi-type process,
$G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}^{d}}\mathbb{P}\left(\boldsymbol{Z}(t)=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}},$
@@
$1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\prod_{i=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}).$ (5.1)
To obtain the moments, we differentiate with respect to $s_{i}$ and evaluate at $\boldsymbol{s}=\boldsymbol{1}$. Note that $01-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})=$ 0. Therefore,
$\ \mathbb{E}\left[Z_{i}(t)\right]=\left.\frac{\partial}{\partial s_{i}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}$ $=\int_{(0,t]}K(x,x)\left.\frac{\partial}{\partial s_{i}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
@@
as the derivatives of the terms in (5.1) for $n>1$ would have a multiple $1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})$ for some $i$, which is zero when evaluated at $\boldsymbol{s}=\boldsymbol{1}$. Similarly for the second moments, only the terms for $n\leq 2$ are relevant, so
$\mathbb{E}\left[Z_{i}(t)Z_{j}(t)\right]=\left.\frac{\partial}{\partial s_{i}\partial s_{j}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}=\int_{(0,t]}K(x,x)\left.\frac{\partial}{\partial s_{i}\partial s_{j}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
$\qquad\quad\,\,\,-\frac{1}{2!}\int_{(0,t]^{2}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq 2}\left.\frac{\partial}{\partial s_{i}\partial s_{j}}\left(\prod_{i=1}^{2}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
@@
$\quad\,\,+\frac{1}{2}\int_{(0,t]}\Big{(}K(x_{1},x_{1})K(x_{2},x_{2})-K^{2}(x_{1},x_{2})\Big{)}\Big{(}2\mathbb{E}\left[Z_{\times,i}(t-x_{1})Z_{\times,j}(t-x_{2})\Big{)}\right]\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2}).$
@@
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp\left(\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle\right)\right],$
and use a similar technique to [35, Theorem 3]. To facilitate the formalism, define $\boldsymbol{Z}(t)=0$ for $t<0$. Using (4.5), we have that for positive $x$,
@@
$\mathcal{L}_{\boldsymbol{Z}_{\times}}\left((t-x),\boldsymbol{s}e^{-\rho t}\right):=\mathbb{E}\left[\exp\left(-\langle\boldsymbol{Z}(t-x),\boldsymbol{s}e^{-\rho t}\rangle\right)\right]\xrightarrow[t\to\infty]{}\mathcal{L}_{\boldsymbol{v}{W}_{\times}}\left(\boldsymbol{s}e^{-\rho x}\right).$ (5.2)
@@
$\mathcal{L}_{\boldsymbol{Z}}\left(t,e^{-\rho t}\boldsymbol{s}\right)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\Lambda(\mathrm{d}\boldsymbol{x}),$ (5.3)
where for readability we introduced the functions, for a vector $\boldsymbol{x}\in\mathbb{R}^{n}$ with $\boldsymbol{x}=(x_{1},\dots,x_{n})$,
$D(\boldsymbol{x}):=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n},\quad\text{and}\quad\Lambda(\mathrm{d}\boldsymbol{x}):=\Lambda(\mathrm{d}x_{1})\cdots\Lambda(\mathrm{d}x_{n}).$
@@
$\mathbbm{1}_{\left\{\boldsymbol{x}\in(0,t]^{n}\right\}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\xrightarrow[t\to\infty]{}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{v}W_{\times}}(\boldsymbol{s}e^{-\rho x})\right).$ (5.4)
To use the last result under the integral sign in (5.3), we need to provide a bound with an integrable function. First, by Hadamard’s inequality, see e.g. [3, (15.A.1)],
$D(\boldsymbol{x})=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\leq\prod_{i=1}^{n}K(x_{i},x_{i}).$ (5.5)
Second, note that as we are working with Laplace transforms of positive random variables, we can assume that the vector $\boldsymbol{s}$ has positive entries. Let $C_{\boldsymbol{s}}:=\max_{j}s_{j}/\min v_{j}$ so that for all $j$, $0\leq s_{j}\leq C_{\boldsymbol{s}}v_{j}$. We next bound as in [35, Section 3.1],
$\begin{split}e^{-\rho t}\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]&=e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]\leq C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{v}\rangle\right]\\ &=C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho\cdot 0}\boldsymbol{Z}_{\times}(0),\boldsymbol{v}\rangle\right]=C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle\boldsymbol{I}_{0},\boldsymbol{v}\rangle\right]=:C_{\boldsymbol{s}}e^{-\rho x},\end{split}$ (5.6)
@@
The last is indeed a $\Lambda(\mathrm{d}\boldsymbol{x})$-integrable because by assumption $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)$ is finite, so
@@
[…]

### v2

@@ 6 Proofs
Proof.
We start with the generating function of the process, similarly to (4.3), the generating function of the multitype process,
$G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}_{0}^{d}}\mathbb{P}\left(\boldsymbol{Z}(t)=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}},$
@@
$1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\Lambda(\mathrm{d}\boldsymbol{x}).$ (6.1)
To obtain the moments, we differentiate with respect to $s_{i}$ and evaluate at $\boldsymbol{s}=\boldsymbol{1}$. Note that we can interchange $\partial/\partial s_{i}$ with the integral, because $G_{Z_{\times}}(t,\boldsymbol{s})\in[0,1]$ for each $t$ and $\boldsymbol{s}||\leq 1$, and the assumption that the expectation of the progeny is finite gives that $\mathbb{E}\left[Z_{\times,i}(t)\right]$ is finite for each $t$, see for example [2, (5) on p.202]. Therefore,
$\left|\frac{\partial}{\partial s_{i}}\prod_{j=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{j},\boldsymbol{s})\right)\right|\leq\sum_{i=1}^{n}\mathbb{E}\left[Z_{\times,i}(t-x_{j})\right]\leq C_{t}$
for some universal $C_{t}\in\mathbb{R}$ as $t$ is fixed. Moreover, as $K$ and $\Lambda$ are locally bounded, we can apply a variation of the dominated convergence theorem, e.g. [20, Theorem 6.28]. To further exchange the summation and $\partial/\partial s_{i}$, by Hadamard’s inequality, see e.g. [3, (15.A.1)],
$D(\boldsymbol{x})=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\leq\prod_{i=1}^{n}K(x_{i},x_{i}),$ (6.2)
so
$\left|\int_{(0,t]^{n}}D(\boldsymbol{x})\frac{\partial}{\partial s_{i}}\prod_{i=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\Lambda(\mathrm{d}\boldsymbol{x})\right|\leq\left(C_{t}\int_{(0,t]}K(x,x)\Lambda(\mathrm{d}x)\right)^{n}=:D_{t}^{n},$
with $D_{t}\in\mathbb{R}_{+}$, and again by dominated convergence, we can interchange $\partial/\partial s_{i}$ and $\sum_{n\geq 1}$ (considering the latter as an integral w.r.t. to a Dirac-type measure).
Observe that $1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})=0$, so
$\ \mathbb{E}\left[Z_{i}(t)\right]=\left.\frac{\partial}{\partial s_{i}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}$ $=-\int_{(0,t]}K(x,x)\left.\frac{\partial}{\partial s_{i}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
@@
as the derivatives of the terms in (6.1) for $n>1$ would have a multiple $1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})$ for some $i$, which is zero when evaluated at $\boldsymbol{s}=\boldsymbol{1}$.
Similarly for the second moments, we can exchange the derivative and sum, because we assume that the respective second moments exist, see (H). We are thus interested in
$1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\left(\left.\frac{\partial^{2}}{\partial s_{i}\partial s_{j}}\prod_{i=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\right)\Lambda(\mathrm{d}\boldsymbol{x}).$ (6.3)
but again since $1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})=0$, the non-zero terms at $\boldsymbol{s}=1$ are obtained for $n=1$ and $n=2$. For $n=1$, we have
$\left.\frac{\partial^{2}}{\partial s_{i}\partial s_{j}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}=-\mathbb{E}\left[Z_{\times,i}(t-x)Z_{\times,j}(t-x)\right],$
and for $n=2$,
${\frac{\partial}{\partial s_{i}}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{1},\boldsymbol{s})\right)$ ${\frac{\partial}{\partial s_{j}}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{2},\boldsymbol{s})\right)$
$+{\frac{\partial}{\partial s_{j}}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{1},\boldsymbol{s})\right){\frac{\partial}{\partial s_{i}}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{2},\boldsymbol{s})\right),$
which evaluated at $\boldsymbol{s}=\boldsymbol{1}$ is exactly
$F^{\ast}(i,j,x_{1},x_{2}):=\mathbb{E}\left[Z_{\times,i}(t-x_{1})\right]\mathbb{E}\left[Z_{\times,j}(t-x_{2})\right]+\mathbb{E}\left[Z_{\times,j}(t-x_{1})\right]\mathbb{E}\left[Z_{\times,i}(t-x_{2})\right].$
Writing $F(i,x_{k}):=\mathbb{E}\left[Z_{\times,i}(t-x_{k})\right]$ and using that
we have
$\int_{(0,t]^{2}}$ $K(x_{1},x_{1})K(x_{2},x_{2})F^{\ast}(i,j,x_{1},x_{2})\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
$=\int_{(0,t]}K(x_{1},x_{1})F(i,x_{1})\Lambda(\mathrm{d}x_{1})\int_{(0,t]}K(x_{2},x_{2})F(j,x_{2})\Lambda(\mathrm{d}x_{2})$
$\qquad\qquad\qquad+\int_{(0,t]}K(x_{1},x_{1})F(j,x_{1})\Lambda(\mathrm{d}x_{1})\int_{(0,t]}K(x_{2},x_{2})F(i,x_{2})\Lambda(\mathrm{d}x_{2})$
$=2\mathbb{E}\left[Z_{i}(t)\right]\mathbb{E}\left[Z_{j}(t)\right].$
Similarly, because by definition $K(x,y)=K(y,x)$, one can also check that
$\int_{(0,t]^{2}}K^{2}(x_{1},x_{2})$ $F^{\ast}(i,j,x_{1},x_{2})\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
$=2\int_{(0,t]^{2}}K^{2}(x_{1},x_{2})\mathbb{E}\left[Z_{\times,i}(t-x_{1})\right]\mathbb{E}\left[Z_{\times,i}(t-x_{2})\right]\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2}).$
Therefore, substituting in (6.3), we get
$\mathbb{E}\left[Z_{i}(t)Z_{j}(t)\right]=\left.\frac{\partial^{2}}{\partial s_{i}\partial s_{j}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}=-\int_{(0,t]}K(x,x)\left.\frac{\partial^{2}}{\partial s_{i}\partial s_{j}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
$\qquad\quad\,\,\,+\frac{1}{2!}\int_{(0,t]^{2}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq 2}\left.\frac{\partial^{2}}{\partial s_{i}\partial s_{j}}\left(\prod_{i=1}^{2}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
@@
$\qquad\quad\,\,\,+\frac{1}{2}\int_{(0,t]^{2}}\Big(K(x_{1},x_{1})K(x_{2},x_{2})-K^{2}(x_{1},x_{2})\Big)F^{\ast}(i,j,x_{1},x_{2})\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
$\qquad\quad\,\,\,+\mathbb{E}\left[Z_{i}(t)\right]\mathbb{E}\left[Z_{j}(t)\right]-\int_{(0,t]^{2}}K^{2}(x_{1},x_{2})\mathbb{E}\left[Z_{\times,i}(t-x_{1})\right]\mathbb{E}\left[Z_{\times,i}(t-x_{2})\right]\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2}).$
@@
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp\left(-\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle\right)\right],$
and use a similar technique to [36, Theorem 3]. To facilitate the formalism, define $\boldsymbol{Z}(t)=0$ for $t<0$. Using (4.5), we have that for positive $x$,
@@
$\mathcal{L}_{\boldsymbol{Z}_{\times}}\left((t-x),\boldsymbol{s}e^{-\rho t}\right):=\mathbb{E}\left[\exp\left(-\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}e^{-\rho t}\rangle\right)\right]\xrightarrow[t\to\infty]{}\mathcal{L}_{\boldsymbol{v}{W}_{\times}}\left(\boldsymbol{s}e^{-\rho x}\right).$ (6.4)
@@
$\mathcal{L}_{\boldsymbol{Z}}\left(t,e^{-\rho t}\boldsymbol{s}\right)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\Lambda(\mathrm{d}\boldsymbol{x}).$ (6.5)
@@
$\mathbbm{1}_{\left\{\boldsymbol{x}\in(0,t]^{n}\right\}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\xrightarrow[t\to\infty]{}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{v}W_{\times}}(\boldsymbol{s}e^{-\rho x_{i}})\right).$ (6.6)
To use the last result under the integral sign in (6.5), we need to provide a bound with an integrable function. First, by Hadamard’s inequality (6.2), $D(\boldsymbol{x})\leq\prod_{i=1}^{n}K(x_{i},x_{i}).$ Second, note that as we are working with Laplace transforms of positive random variables, we can assume that the vector $\boldsymbol{s}$ has positive entries. Let $C_{\boldsymbol{s}}^{\ast}:=\max_{j}s_{j}/\min_{j}v_{j}$ so that for all $j$, $0\leq s_{j}\leq C_{\boldsymbol{s}}^{\ast}v_{j}$. We next bound as in [36, Section 3.1],
$\begin{split}&e^{-\rho t}\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]=e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]\\ &\qquad\qquad\leq C_{\boldsymbol{s}}^{\ast}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{v}\rangle\right]\\ &\qquad\qquad=C_{\boldsymbol{s}}^{\ast}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho\cdot 0}\boldsymbol{Z}_{\times}(0),\boldsymbol{v}\rangle\right]=C_{\boldsymbol{s}}^{\ast}e^{-\rho x}\mathbb{E}\left[\langle\boldsymbol{I}_{0},\boldsymbol{v}\rangle\right]=:C_{\boldsymbol{s}}e^{-\rho x},\end{split}$ (6.7)
@@
The last is indeed a $\Lambda(\mathrm{d}\boldsymbol{x})$-integrable function because $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)$ is finite by assumption, so
@@
$\hskip 56.9055pt\xrightarrow[t\to\infty]{}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{v}W_{\times}}(\boldsymbol{s}e^{-\rho x_{i}})\right)\Lambda(\mathrm{d}\boldsymbol{x}).$
@@
Proof of Lemma 5.5.
We will use the following result.
Lemma 6.1 (Lemmas 2 and 3 from [44]).
Let $\alpha_{\infty}$ and $\beta_{\infty}$ be positive real constants. Then we have the following.
  1. 1.
    If $\alpha(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\alpha_{\infty}e^{\delta t}$ with $\delta\in\mathbb{R}$ and $\int_{0}^{\infty}e^{-\delta x}\beta(x)\,\mathrm{d}x$ is finite, then
    $\int_{0}^{t}\beta(t-x)\alpha(x)\mathrm{d}x\stackrel{{\scriptstyle\infty}}{{\sim}}\alpha(t)\int_{0}^{\infty}e^{-\delta x}\beta(x)\mathrm{d}x.$
  2. 2.
    If $\alpha(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\alpha_{\infty}$ and $\beta(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\beta_{\infty}$, then
    $\int_{0}^{t}\beta(t-x)\alpha(x)\mathrm{d}x\stackrel{{\scriptstyle\infty}}{{\sim}}\alpha_{\infty}\beta_{\infty}t.$
Let us start with the observation that the stationarity assumption (5.2), applied for $n=1$, implies that $K(x,x)$ is constant over $x$. Denote its value by $K_{\ast}$. By the classical result, we have see e.g. [2, p. 203] or [41, p.151],
$\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}a_{i}:=e^{\rho t}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i},$ (6.10)
Next, by Theorem 5.1, $\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}\int_{0}^{t}\mathbb{E}\left[Z_{\times,i}(t-x)\right]\lambda(x)\mathrm{d}x$, so
  • if $\delta>\rho$, by (6.10) the integral $\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x$ is finite, and from the first part of Lemma 6.1 with $\alpha(t)=\lambda(t)$ and $\beta(t)=\mathbb{E}\left[Z_{\times,i}(t)\right]$, we obtain
    $\mathbb{E}\left[Z_{i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\delta t}A_{i}:=e^{\delta t}K_{\ast}\lambda_{\infty}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x.$
  • if $\delta=\rho$, we apply the second part of Lemma 6.1 with $\beta(t)=e^{-\delta t}\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}a_{i}$, and $\alpha(t)=e^{-\delta t}\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}$, so
    $\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}e^{\delta t}\int_{0}^{t}\beta(t-x)\alpha(x)\mathrm{d}x\stackrel{{\scriptstyle\infty}}{{\sim}}K_{\ast}e^{\delta t}ta_{i}\lambda_{\infty}=:A_{i}^{\prime}te^{\delta t}.$
We continue with a similar strategy for the variance: the asymptotics of the second moment of $Z_{\times,i}(t)$ are available in [41, p.152] or [2, p.204], giving us, for some positive constants $B_{j,i}$,
$\mathbb{E}\left[\left(Z_{\times,i}(t)\right)^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}\begin{cases}e^{\rho t}B_{1,i}&\text{if $\rho<0$}\\ tB_{2,i}&\text{if $\rho=0$}\\ e^{2\rho t}B_{3,i}&\text{if $\rho>0$}.\end{cases}$
Further, by the second part of Theorem 5.1,
$\operatorname{Var}(Z_{i}(t))\leq K_{\ast}\int_{0}^{t}\mathbb{E}\left[\left(Z_{\times,i}(t-x)\right)^{2}\right]\lambda(x)\mathrm{d}x.$
We consider the relevant cases.
  • Fix $\delta>\rho$. We apply Lemma 6.1 with $\alpha(t)=\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]$ and $\beta(t)=\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\delta t}\lambda_{\infty}$. Note that if $\rho<0$, $\alpha(t)\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}B_{1,i}$, and if $\rho=0$, $\alpha(t)\stackrel{{\scriptstyle\infty}}{{\sim}}tB_{2,i}$, so in both cases $\int_{0}^{\infty}e^{-\delta x}\alpha(x)\mathrm{d}x$ is finite, so we obtain
    $K_{\ast}\int_{0}^{t}\mathbb{E}\left[\left(Z_{\times,i}(t-x)\right)^{2}\right]\lambda(x)\mathrm{d}x\stackrel{{\scriptstyle\infty}}{{\sim}}K_{\ast}e^{\delta t}\lambda_{\infty}\int_{0}^{\infty}e^{-\delta x}\beta(x)\mathrm{d}x=\mathrm{O}\left(e^{\delta t}\right).$
  • Fix $\rho>0$ and $\delta=2\rho$. We apply the second item of Lemma 6.1 with $\alpha(t)=e^{-2\rho t}\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}B_{3,i}$ and $\beta(t)=e^{-\delta t}\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}$, so
    $K_{\ast}\int_{0}^{t}\mathbb{E}\left[\left(Z_{\times,i}(t-x)\right)^{2}\right]\lambda(x)\mathrm{d}x=K_{\ast}e^{2\rho t}\int_{0}^{t}\alpha(t-x)\beta(t)\mathrm{d}t=\mathrm{O}\left(te^{2\rho t}\right)=\mathrm{O}\left(te^{\delta t}\right).$
  • With similar arguments, if $\delta>2\rho>0$, we have $\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{2\rho t}B_{3,i}$, so we apply Lemma 6.1 with $\alpha(t):=\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]$ and $\beta(t)=\lambda(t)$. If $\delta<2\rho$, we should reverse the roles of $\alpha$ and $\beta$.
  • If $\delta=\rho>0$, we use Lemma 5.5 with $\alpha(t)=\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}e^{\delta x}$ and $\beta(t)=\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{2\rho t}B_{3,i}.$